It takes 365.25 days for the earth to complete one revolution around the sun. Calculate the


A. Period


B. Frequency


C. Angular frequency of earth as it revolves around the sun.

1
Expert's answer
2022-01-20T13:50:20-0500

A. The period

"T=365.25\:\rm days=3.1558*10^7\: s"

B. Frequency

"f=1/T=1/3.1558*10^7=3.1688*10^{-8}\:\rm Hz"

C. Angular frequency

"\omega=2\pi f=2\pi*3.1688*10^{-8}\\=1.990*10^{-7}\:\rm rad/s"